OpenStudy (anonymous):

A swimming pool is partially drained using a submersible pump. The volume of water in the pool changes steadily by negative eight hundred gallons over twentyminutes. What is the change in volume during each minute? negative one thousand six hundred gallons negative forty gallons forty gallons one thousand six hundred gallons

OpenStudy (anonymous):

i got it! -40

OpenStudy (anonymous):

Is it correct?

OpenStudy (tkhunny):

-800/20 = -80/2 = -40 -- Why do you doubt?